# DPNet: Doppler LiDAR Motion Planning for Highly-Dynamic Environments

### v1

@@ IV-A2 Kalman Gain Learning
where $\hat{\mathbf{x}}^{n}_{h+1|h}$ is the priori state prediction and $\hat{\mathbf{y}}^{n}_{h+1|h}$ is the estimated state observation using forward propagation. After $\Delta t$, LiDAR scan $\mathcal{P}_{h+1}$ arrives, providing the actual $\mathbf{y}^{n}_{h+1}$ by equations (3) and (6). We then update the posteriori state prediction by
@@
Particularly, we adopt an RNN [revach2022kalmannet] to learn a robust and adapative $\mathcal{K}_{h+1}^{n}$ for posteriori update, thereby alleviating the noises originated from either mismatched $\mathbf{T}$ or inaccurate ${\mathbf{y}}_{h+1|h}^{n}$. The entire procedure is summarized in Algorithm 1, where $\hat{\mathcal{O}}_{t}^{n}=\{\hat{\mathbb{O}}_{h+1|t}^{n}(\hat{\mathbf{x}}_{h+1|t}^{n})|\forall h\in\mathcal{H}_{t}\}$ represents the set of bounding box predictions over the planning horizon.
@@ IV-B Doppler-Tuned MPC for Planning
With the Doppler-learned $\{\hat{\mathcal{O}}_{0}^{n},\hat{\mathcal{O}}_{1}^{n},\cdots\}_{n=1}^{N}$ from D-KalmanNet (Algorithm 1), we propose DT-MPC which enables real-time controller tuning by leveraging Doppler-inferred collision check. According to [cheng2024difftune, tao2024difftune], the core of controller tuning is transforming the fixed value $d_{\text{safe}}$ in $\mathsf{P}_{t}$ into a learnable parameter vector $\bm{\phi}=[\phi_{t+1,1},\cdots,\phi_{t+H,N}]^{T}$. As such, $\mathsf{P}_{t}$ becomes a parameter optimization problem
@@
where $\mathcal{L}$ is the closed-loop collision cost function, $\gamma$ is the collision penalty coefficient and $\Phi$ defines the upper and lower bounds of $\bm{\phi}$. One example of $\mathcal{L}$ is the quadratic loss of the collision penalty [han2025neupan], where
@@ IV-B1 Doppler-Inferred Collision Check
In conventional auto-tuning MPC [han2025neupan], $\mathsf{Q}_{t}$ is solved by differentiable optimization with gradient update to yield a larger $\bm{\phi}$ when the robot collides with any obstacle, yet can be inefficient if collision events are sparse. In contrast, in our system, we have highly-accurate (as demonstrated in Section V-B) prediction $\hat{\mathbb{O}}_{h+1|t}^{n}$ provided by D-KalmanNet. Moreover, for an MPC system, although solving $\mathsf{Q}_{t}$ provides the complete solution sequence of $\mathcal{W}_{t}$ and $\mathcal{S}_{t}$ along $\mathcal{H}_{t}$, only the first action $\mathbf{w}_{t}$ is executed and ends up with state $\mathbf{s}_{t+1}$, resulting in a sequence of non-executed actions $\mathcal{W}_{t}^{\circ}=\{\mathbf{w}_{h}\}_{h=t+1}^{t+H-1}$ and a sequence of potential states $\mathcal{S}_{t}^{\circ}=\{\mathbf{s}_{h}\}_{h=t+2}^{t+H}$. Hence, by leveraging $\mathcal{W}_{t}^{\circ}$ and $\mathcal{S}_{t}^{\circ}$ as a prediction of ego-trajectory, and obstacle-trajectory prediction $\hat{\mathbb{O}}_{h+1|t}^{n}$ from D-KalmanNet, a Doppler-inferred collision check can be formulated to enable run-time MPC tuning. Specifically, let $\mathcal{H}_{t-1}^{-}=\mathcal{H}_{t-1}\setminus\{t-1\}$ denote the truncated historical horizon at time $t$. The bounding box sequence of historical potential states $\mathcal{S}_{t-1}^{\circ}$ is denoted as
$\mathcal{G}_{t-1}^{\circ}=\{\mathbb{G}^{\circ}_{h+1}(\mathbf{s}_{h+1})|\mathbf{s}_{h+1}\in\mathcal{S}_{t-1}^{\circ},\forall h\in\mathcal{H}_{t-1}^{-}\},$ (12)
Since $\mathcal{H}_{t-1}^{-}$ ends at t+H2t+H-2, the last element of $\hat{\mathcal{O}}_{t}^{n}$ is also truncated to match the length. The criteria of potential collision along the truncated historical horizon is
@@
where $d_{0}$ is the collision threshold.
Figure 4: Qualitative analysis of robot motions (in ROS-Rviz and Carla [dosovitskiy2017carla] views) and corresponding control commands. Static obstacles are marked as green boxes, while dynamic ones are red boxes with arrows indicating their moving directions. The blue line is the path connecting start and goal points. The dots in front of dynamic obstacles represent motion predictions.
@@ IV-B2 Run-Time Auto-Tuning
Now, if Doppler-inferred collision (13) exists for any obstacle, the MPC parameters $\bm{\phi}$ for current planning round can be auto-tuned to cope with potential future risks. Specifically, $\phi_{h,n}$ is decomposed into a constant component and an adaptive component as $\phi_{h,n}=d_{1}+\tau(n,h)d_{2}$, where $d_{1}$ is the compulsory minimum distance and $\tau(n,h)d_{2}$ is the Doppler adaptive distance. This converts tuning $\phi_{h,n}$ to tuning the factor $\tau(n,h)$, and allows us to adopt a simple yet effective parameter update method to surrogate for the gradient $\partial\mathcal{L}/\partial\tau(n,h)$. Denoting $\tau(n,h)=\tau_{1}(n)\tau_{2}(h)$, where $\tau_{1}(n)\in(0,1)$ is the spatial factor and $\tau_{2}(h)\in(0,1)$ is the temporal factor, the update is performed as follows:
First, the spatial factor $\tau_{1}(n)$ measures the imminence of the Doppler-inferred collision by
@@
where $\tau_{2;\min}$ is the lower bound and $\beta\in(0,1)$ is the horizon decay rate. A larger $\beta$ imposes stricter constraints on nearby predictions while relaxing constraints for distant ones.
@@ IV-B3 Solution by ADMM
The only nonconvex part in $\mathsf{R}_{t}$ is $\mathcal{L}(\hat{\bm{\phi}},\{\hat{\mathcal{O}}_{t}^{n}\})$. To tackle this function, we can equivalently transform $\mathcal{L}$ into a biconvex form using Lagrange duality [han2023rda, han2025neupan], and then solve the resultant biconvex version by ADMM. This ADMM will converge to a stationary point of $\mathsf{R}_{t}$.
@@ V Experiments
We implement DPNet in ROS Noetic on Linux system. We compare DPNet with state-of-the-art methods and provide both qualitative and quantitative results in heterogeneous scenarios. The experiments are conducted using Carla simulator with Doppler LiDAR integration [carla-aeva]. Moreover, we evaluate the proposed D-KalmanNet against existing obstacle tracking methods to assess the robustness in highly-dynamic environments and on resource-constrained platforms. The D-KalmanNet is trained for 2000 epochs using Doppler LiDAR measurements and ground truth vehicle trajectories, using the real-world dataset AevaScenes [aevascenes].
@@ V-A End-to-End Evaluation in Carla
We compare the following methods [^1] Reinforcement learning baselines are not included, since they suffer from a lack of safety assurance in highly-dynamic scenarios. For example, we have tested twin delayed deep deterministic policy gradient algorithm (TD3) in IR-SIM [han2023rda], with $15$ obstacles randomly moving at $2$ m/s in a $40\times 40m$ square. The success rate of TD3 is only 70/100. for end-to-end evaluations in Carla-ROS simulation: (1) DPNet: our proposed method; (2) MPC-D-CBF [jian2023dynamic]: an MPC-based planner that predicts obstacle motion using LiDAR and Kalman filter, constructing D-CBFs online for navigation among moving obstacles; (3) RDA [han2023rda]: an MPC-based planner that uses parallelizable dual ADMM to accelerate agile collision avoidance; (4) OBCA [zhang2020optimization]: an MPC-based planner with optimization-based collision avoidance for convex obstacles; (5) DPNet (Ablation): our proposed method where DT-MPC is disabled. All methods are set up with $H=15$ and $\Delta t=0.12$s. Ego robot and obstacles are modeled as car-like wheel robots.
@@
[zhang2020optimization] [han2023rda] [jian2023dynamic] (Ablation)
- 1D-1S clutter level
AvgAcc $\downarrow$ 1.189 1.196 1.186 $\cellcolorpurple21.17$0
[…]
- 2D-2S clutter level
[…]
MaxAcc $\downarrow$ 7.744 8.689 8.036 $\cellcolorpurple27.01$9
[…]
- 4D-4S clutter level
@@
TABLE I: Performance comparison at different clutter levels. Purple denotes the best performance.
Qualitative Analysis. As shown in Fig. 4, this scenario consists of two static obstacles and two rapid-motion obstacles (at 55\,m/s), one with vertical direction and the other with lateral direction. These obstacles will not avoid the ego robot. The ego robot needs to traverse the area while avoiding collisions with any obstacle. It can be seen that our proposed DPNet succeeds to pass through the area with agile motions. Particularly, at t=2t=2\,s, if velocity information is ignored, directly passing between the two dynamic objects appears to be the optimal choice. This is case of RDA, which crashes into the lateral moving obstacle at t=5.6t=5.6\,s. The is because at t=4t=4\,s, the ego-robot is completely surrounded by four obstacles, and it was too late to steer by then. In contrast, with accurate velocity understanding, our DPNet can look ahead into the future, “imaging” that directly passing through is risky, and thus early-executes the left-turn at t=2t=2\,s. As such, at t=3.6t=3.6\,s and t=5.2t=5.2\,s, the ego-robot bypasses from behind the laterally moving obstacle. This demonstrates the power of Doppler model-based learning in handling highly-dynamic environments. In terms of other methods, the OBCA planner crashes into the vertical moving obstacle due to its delayed motions caused by low computation frequency. The MPC-D-CBF planner makes over-conservative decision upon handling the vertical moving obstacle due to its inaccurate velocity understandings, resulting in a crash finally.
Quantitative Analysis. As shown in Fig. 6, we set up 4 scenes with different environmental clutter levels to evaluate detailed performances. A higher clutter level means more dynamic and static obstacles, with dynamic obstacles assigned fixed directions and a velocity of 55\,m/s. The following six metrics are used: (a) AvgAcc (m/s2): the average acceleration during the execution; (b) MaxAcc (m/s2): the maximum acceleration duing the execution; (c) AvgJerk (m/s3): the average jerk during the execution; (d) IteTime (s): the average iteration time of the solver during the execution; (e) PassTime (s): the average time duration to traverse the area; (f) PassRate (%): the rate of successfully traversing the area without collisions. Experiments are executed $100$ times for each setting.
Results are reported in Table I. As clutter level increases, DPNet consistently outperforms all benchmark methods, in terms of both motion smoothness and time efficiency. Furthermore, ablation studies demonstrate that DT-MPC effectively improves the navigation performance compared to its non-DT counterpart, at the cost of a slight increase in execution time. This confirms the enhanced adaptability of DT-MPC to properly tune safety distances online, especially in high clutter levels. To further assess DPNet’s agility in highly-dynamic environments, we compare pass rate under varying obstacle velocity conditions. Experiments are conducted in the 2D-2S scene with three different obstacle velocities: 5 m/s, 7.5 m/s, and 10 m/s. As illustrated in Fig. 6, results demonstrate that DPNet maintains a high pass rate across all velocities, highlighting its robustness in the presence of fast-moving obstacles.
Figure 5: Clutter levels. D (S) refers to dynamic (static) obstacles.
Figure 6: Pass rate comparison at different obstacle velocities in 2D-2S clutter level.
Figure 7: Visualization of critical predictions for a rapidly approaching vehicle (at about 25 m/s, highlighted in orange) in Highway scenario of AevaScene [aevascenes]. The upper half shows Doppler LiDAR’s environmental dynamics detection (outgoing / incoming bounding boxes marked in red / blue). The lower half provides detailed comparison of trajectory predictions for the threat.
@@ V-B Evaluation on Real-World Datasets
To evaluate D-KalmanNet’s robustness to real-world sensor noises, we conduct experiments on the AevaScene [aevascenes] dataset, which contains on-road Doppler LiDAR data and ground truth annotations of vehicle trajectories in both Highway and City scenarios. Following methods are compared: (1) D-KNet: our proposed tracking method; (2) KNet [revach2022kalmannet]: the KalmanNet, a physics-informed deep learning method; (3) KF [jian2023dynamic]: the Kalman filter, a widely-used conventional tracking method; (4) D-KF [peng2021detection]: the conventional Kalman filter aided by Doppler LiDAR.
To quantify tracking accuracy, we use the normalized mean squared error (NMSE) in decibels (dB) following [revach2022kalmannet]. Specifically, for each vehicle, and at every feasible time step $t$ , we first evaluate the horizon-averaged NMSE over $\mathcal{H}_{t}$:

### v2

@@ IV-A2 Kalman Gain Learning
where $\hat{\mathbf{x}}^{n}_{h+1|h}$ is the priori state prediction and $\hat{\mathbf{y}}^{n}_{h+1|h}$ is the estimated state observation using forward propagation. After $\Delta t$, new scan $\mathcal{P}_{h+1}$ arrives, providing the actual $\mathbf{y}^{n}_{h+1}$ by (2) and (4). We then update the posteriori state prediction by
@@
We adopt an RNN [17] to learn a robust and adapative $\mathcal{K}_{h+1}^{n}$ in (7) for posteriori update, thereby alleviating the noises originated from either mismatched $\mathbf{T}$ or inaccurate ${\mathbf{y}}_{h+1|h}^{n}$. The entire procedure is summarized in Algorithm 1, where $\hat{\mathcal{O}}_{t}^{n}=\{\hat{\mathbb{O}}_{h+1|t}^{n}(\hat{\mathbf{x}}_{h+1|t}^{n})\}_{\forall h\in\mathcal{H}_{t}}$ represents bounding box predictions over the planning horizon.
Algorithm 2 Doppler Collision Check-and-Tune
1: Initialize $\mathcal{S}_{t-1}=\emptyset$
2:while True do
3:  $\varrho_{n}\leftarrow\infty$ for $n\in\mathcal{N}$, $\kappa\leftarrow\kappa^{\text{init}}$   // New Iteration
4:  Obtain $\hat{\mathcal{O}}_{t}^{n}$ from the D-KalmanNet
5:  for $h\in\mathcal{H}_{t-1}^{-}$ do
6:   for $n\in\mathcal{N}$ do
7:    if $\varrho_{n}=\infty$ and $\mathbf{dist}(\mathbb{G}^{\circ}_{h+1},\hat{\mathbb{O}}_{h+1|t}^{n})\leq d_{0}$ then
8:     $\varrho_{n}\leftarrow\kappa$   // Assign Collision Priority
9:    end if
10:   end for
11:   Update $\kappa\leftarrow\kappa+\Delta\kappa$   // Update Accumulator
12:  end for
13:  $t\leftarrow t+1$
14:end while
@@ IV-B Doppler-Tuned MPC for Planning
A fixed $d_{\text{safe}}$ in $\mathsf{P}_{t}$ often leads to poor performance or problem infeasibility [22]. This can be addressed by MPC auto-tuning [6], which transforms $d_{\text{safe}}$ into a learnable parameter vector $\bm{\phi}=[\phi_{t+1,1},\cdots,\phi_{t+H,N}]^{T}$ and reformulates $\mathsf{P}_{t}$ into the following parameter optimization problem:
@@
where $\mathcal{L}$ is the closed-loop collision cost function, $\gamma$ is the collision penalty coefficient and $\Phi$ defines the upper and lower bounds of $\bm{\phi}$. One example of $\mathcal{L}$ is the quadratic loss of the collision penalty [6], where
@@
To minimize $\mathcal{L}$, DT-MPC adopts a heuristic rule [26] to trigger parameter updates by using $\hat{\mathcal{O}}^{n}_{t}$ for Doppler-inferred collision check. This allows exploring whether real-time Doppler measurements can facilitate tuning even with simple heuristics (see Section V-D for its limitation).
@@ IV-B1 Doppler-Inferred Collision Check
In an MPC system, solving $\mathsf{Q}_{t}$ produces complete $\mathcal{W}_{t}$ and $\mathcal{S}_{t}$ along $\mathcal{H}_{t}$, while only the first action $\mathbf{w}_{t}$ is executed, which ends up with state $\mathbf{s}_{t+1}$. This leaves the potential state sequence $\mathcal{S}_{t}^{\circ}=\{\mathbf{s}_{h}^{\circ}\}_{h=t+2}^{t+H}$ as a natural surrogate of ego trajectory prediction. Hence, before solving $\mathsf{Q}_{t}$, we perform parameter tuning by assessing historical potential state sequence $\mathcal{S}_{t-1}^{\circ}$ from the solution of $\mathsf{Q}_{t-1}$. Specifically, $\mathcal{S}_{t-1}^{\circ}$ corresponds to the ego bounding box prediction $\{\mathbb{G}^{\circ}_{h+1}(\mathbf{s}_{h+1}^{\circ})\}_{\forall h\in\mathcal{H}_{t-1}^{-}}$, where $\mathcal{H}_{t-1}^{-}=\mathcal{H}_{t-1}\setminus\{t-1\}$ denotes the truncated historical horizon. Inspired by [18], we then check potential collisions between $\{\hat{\mathbb{O}}_{h+1|t}^{n}(\hat{\mathbf{x}}_{h+1|t}^{n})\}$ and $\{\mathbb{G}^{\circ}_{h+1}(\mathbf{s}_{h+1}^{\circ})\}$ along $\mathcal{H}_{t-1}^{-}$} , using the following criteria:
@@
where $d_{0}$ is the distance threshold of Doppler-inferred collision. Note that the last element of $\hat{\mathcal{O}}_{t}^{n}$ is excluded from the check to match the length of $\mathcal{H}_{t-1}^{-}$.
Figure 4: Qualitative analysis of robot motions (in ROS-RViz and Carla [3] views) and corresponding control commands. Static obstacles are marked as green boxes, while dynamic ones are red boxes with arrows indicating their moving directions. The blue line is the path connecting start and goal points. The dots in front of dynamic obstacles represent motion predictions.
@@ IV-B2 Run-Time Auto-Tuning
If such Doppler-inferred collision (10) exists for any obstacle, $\bm{\phi}$ for $\mathsf{Q}_{t}$ is updated to alleviate potential future risks. We first decompose $\phi_{h,n}$ into a constant component and an adaptive component as $\phi_{h,n}=d_{1}+\tau(n,h)d_{2}$, where $d_{1}$ is the compulsory minimum distance and $\tau(n,h)d_{2}$ is the Doppler adaptive distance. This converts tuning $\phi_{h,n}$ to tuning the factor $\tau(n,h)$. We further decompose $\tau(n,h)=\tau_{1}(n)\tau_{2}(h)$, where $\tau_{1}(n)\in(0,1)$ is the spatial factor and $\tau_{2}(h)\in(0,1)$ is the temporal factor. First, the spatial factor $\tau_{1}(n)$ measures the imminence of the Doppler-inferred collision by
@@
where $\tau_{2;\min}$ is the lower bound and $\beta\in(0,1)$ is the horizon decay rate. A larger $\beta$ imposes stricter constraints on nearby predictions while relaxing constraints for distant ones. Table II in Section V provides a sensitivity analysis of $(\alpha,\beta)$.
@@ V Experiments
We implement DPNet in ROS Noetic on Linux system, using Carla simulator with Doppler LiDAR integration [1]. We compare DPNet with state-of-the-art methods and provide both qualitative and quantitative results in highly-dynamic environments randomize by DynaBARN [14] benchmark, along with ablation study and sensitivity analysis for DT-MPC. We also evaluate D-KalmanNet against existing obstacle tracking methods in highly-dynamic environments, and assess its efficiency on resource-constrained platform NVIDIA Jetson Orin NX. Ground truths of obstacle bounding boxes are used to group Doppler LiDAR points.
We train D-KalmanNet for 2000 epochs on the real-world Doppler LiDAR dataset AevaScenes [15] comprising 100 sequences (50 in City, 50 in Highway). Each sequence contains 100 frames recorded at 10 Hz with ground-truth vehicle linear velocities and bounding box poses. Training uses 80 randomly selected sequences (40 per scenario), and the remaining 20 sequences serve as the evaluation set in Section V-B.
@@ V-A End-to-End Evaluation in Carla
We compare the following methods [^1] Reinforcement learning baselines are not included, since they suffer from a lack of safety assurance in highly-dynamic scenarios. for end-to-end evaluations in Carla-ROS simulation: (1) DPNet: our proposed method; (2) MPC-D-CBF [10]: an MPC-based planner that predicts obstacle motion using LiDAR and Kalman filter, constructing D-CBFs online for navigation among moving obstacles; (3) RDA [7]: an MPC-based planner that uses parallelizable dual ADMM to accelerate collision avoidance; (4) OBCA [27]: an MPC-based planner with optimization-based collision avoidance for convex obstacles; (5) DPNet (Ablation): our proposed method where DT-MPC is disabled. All methods are set up with $H=15$ and $\Delta t=0.1\,$s. DPNet’s specific parameters are set as follows: $d_{0}=0.5\,$m, $d_{1}=0.1\,$m, $d_{2}=0.4\,$m, $\alpha=0.2$, $\beta=0.05$, $\tau_{1;\min}=0.3$, $\tau_{2;\min}=0.3$, $\kappa^{\text{init}}=1.0$, and $\Delta\kappa=1.0$. Ego robot and obstacles are modeled as car-like wheel robots.
Qualitative Analysis. As shown in Fig. 4, this scenario consists of two static obstacles and two rapid-motion obstacles (at 55\,m/s) with frontal and lateral directions. These obstacles will not avoid the ego robot. The ego robot needs to traverse the area without any collisions. It can be seen that our DPNet succeeds to pass through the area with agile motions. Particularly, at t=2t=2\,s, if velocity information is ignored, directly passing between the two dynamic objects appears to be the optimal choice. However, at t=4t=4\,s, the ego-robot is completely surrounded by four obstacles, and it was too late to steer by then. This is case of RDA, which crashes into the lateral moving obstacle at t=5.6t=5.6\,s. In contrast, with accurate velocity understanding, our DPNet can look ahead into the future, “imagining” that directly passing through is risky, and thus early-executes the left-turn at t=2t=2\,s. As such, at t=3.6t=3.6\,s and t=5.2t=5.2\,s, the ego-robot bypasses from behind the laterally moving obstacle. This demonstrates the power of Doppler model-based learning in handling highly-dynamic environments. In terms of other methods, OBCA crashes into the frontally moving obstacle due to its delayed motions caused by low computation frequency. MPC-D-CBF planner makes over-conservative decision upon handling the frontally moving obstacle due to its inaccurate velocity understandings, resulting in a crash eventually.
Quantitative Analysis. To evaluate DPNet in highly-dynamic environments, we adopt DynaBARN [14] benchmark to set up obstacles with randomized trajectories, velocities ($6\pm 2\,$m/s), and accelerations. The following metrics are used: (a) AvgAcc (m/s2): the average acceleration during the execution; (b) MaxAcc (m/s2): the maximum acceleration duing the execution; (c) AvgJerk (m/s3): the average jerk during the execution; (d) IteTime (ms): the average iteration time of the solver during the execution; (e) PassTime (s): the average time duration to traverse the area; (f) PassRate (%): the rate of successfully traversing the area without collisions.
Results are reported in Table I, where each obstacle setting is executed 100 times. As the number of obstacles varies from 1 to 7, DPNet consistently outperforms other baselines, with a significantly superior PassRate in dense obstacle settings. In terms of computational efficiency, DPNet achieves a low IteTime, close to that of RDA [7]. However, PassRate of RDA is far below that of DPNet, e.g., degrading by $42.2\,\%$ when the number of obstacles is 7. These results confirm that DPNet robustifies collision avoidance with efficient computation. Ablation study shows that without DT-MPC, DPNet’s PassRate drops especially under high obstacle numbers, demonstrating the effectiveness of leveraging real-time Doppler measurements for MPC tuning. Fig. 5 demonstrates how DPNet traverses an area with 5 randomized highly-dynamic obstacles.
Table II presents a sensitivity analysis of DT-MPC under different parameter settings of ($\alpha$, $\beta$), with 100 runs for each setting. DT-MPC maintains stable performance across various settings. Increasing $\alpha$ or $\beta$ slightly improves PassRate while degrading AvgJerk, since larger $\alpha$ or $\beta$ leads to more conservative actions (e.g., abrupt turns) for closer obstacles. Note that DT-MPC achieves a low tuning time $t_{\text{DT-MPC}}$ (<1.5<1.5\,ms), demonstrating its real-time efficiency.
TABLE I: Quantitative Comparison
@@
[27] [7] [10] (Ablation)
- 1 obstacle
@@
- 5 obstacles
[…]
- 7 obstacles
[…]
@@ V-B Evaluation on Real-World Datasets
To evaluate D-KalmanNet’s robustness to real-world sensor noises, we conduct experiments on AevaScene [15] dataset and compare the following methods: (1) D-KNet: our proposed D-KalmanNet; (2) KNet [17]: the KalmanNet, a physics-informed deep learning method; (3) KF [10]: the Kalman filter, a widely-used conventional tracking method; (4) D-KF [16]: the conventional Kalman filter aided by Doppler LiDAR.
To quantify tracking accuracy, we use the normalized mean squared error (NMSE) in decibels (dB) following [17]. Specifically, for each vehicle, and at every feasible time step $t$ , we first evaluate the horizon-averaged NMSE over $\mathcal{H}_{t}$:
@@
TABLE II: Sensitivity Analysis of DT-MPC
Parameter 3 obstacles 5 obstacles 7 obstacles
[…]
default 8.058 88.0 9.134 57.0 9.056 45.0
[3pt/3pt] $\alpha=1.0$ 8.264 92.0 9.001 59.0 9.077 44.0
$\alpha=0.5$ 8.179 89.0 8.972 54.0 8.796 47.0
$\alpha=0.1$ 7.953 86.0 8.799 48.0 8.436 40.0
[3pt/3pt] $\beta=1.0$ 8.349 90.0 9.206 61.0 9.165 48.0
$\beta=0.5$ 8.370 79.0 8.715 55.0 9.033 39.0
$\beta=0.1$ 8.169 81.0 8.984 57.0 8.912 43.0
$\bm{t}_{\text{DT-MPC}}\,(\text{ms})$ 0.621 0.997 1.498
Figure 5: Visualization of DPNet traversing a highly-dynamic 5-obstacle area, randomized by DynaBARN [14] benchmark. Frames are sampled by $1.25$ second interval.
Qualitative Analysis. Fig. 6 demonstrates a visualized result of tracking one rapidly approaching vehicle (at about 25 m/s) on the Highway scenario. Among the evaluated methods, D-KNet demonstrates superior performance by effectively leveraging Doppler cues to enable accurate, real-time tracking of rapid objects, achieving the lowest tracking error of 21.18–21.18 dB. In contrast, KF and KNet struggle to handle the approaching vehicle with abrupt accelerations. This comparison highlights the importance of integrating model-based learning and Doppler LiDAR measurements for more robust tracking under real-world settings.
Figure 6: Visualization of critical predictions for a rapidly approaching vehicle (at about 25 m/s, highlighted in orange) in Highway scenario of AevaScene [15]. The upper half shows Doppler LiDAR’s environmental dynamics detection (outgoing / incoming bounding boxes marked in red / blue). The lower half provides detailed comparison of trajectory predictions for the threat.
TABLE III: Prediction NMSE (dB) by Mean $\pm$ Std